# Lattice dynamics study of electron-correlation-induced charge density wave in antiferromagnetic kagome metal FeGe

### v1

@@ III Summary
Starting from the initial P6/mmm structure, which contains an ideal Fe kagome net, we found that there exists a critical value of the Hubbard parameter (around eV), for which the softened phonon mode becomes imaginary. From analysis of the most predominant imaginary soft mode at the L point, we found a new structure with the Ibam symmetry. This structure comes from P6/mmm due to the shift of Ge atoms within the Fe kagome net layer. However, the lattice dynamics of Ibam structure reveals that the soft mode with imaginary frequency still exists at the point. In this case, the displacement introduced by this soft mode associated with the shift of Ge atoms in a direction perpendicular to the Fe kagome net leads to a stable Immm structure.
@@
Additionally, we performed calculations for the slab structure with the Immm symmetry to generate STM images for the Fe-kagome-like and Ge-deformed honeycomb surface terminations. The calculated STM topographies indicate the existence of charge ordering. The observed CDW phase is associated with a deformed Ge honeycomb-like lattice.
Appendix A Computational techniques
The first-principles density functional theory (DFT) calculations were performed using the projector augmented-wave (PAW) potentials [87] implemented in the Vienna Ab initio Simulation Package (Vasp) code [88, 89, 90]. For the exchange-correlation energy, the generalized gradient approximation (GGA) in the Perdew–Burke–Ernzerhof (PBE) parametrization was used [91]. The energy cutoff for the plane-wave expansion was set to eV. We introduced the correlation effect on Fe orbitals using the DFT+U scheme proposed by Dudarev et al. [92].
The optimization of the lattice constants and atomic positions in the presence of the spin–orbit coupling (SOC) was performed for magnetic unit cells (with AFM order). The structures were optimized with different k-grids generated using the Monkhorst–Pack scheme [93] depending on the investigated symmetry. The initial structure with the P6/mmm symmetry was optimized using k–point grid. Structures with the Ibam and Immm symmetries were optimized with k–point grids. As a convergence criterion of the optimization loop, we took the energy change below eV and eV for ionic and electronic degrees of freedom, respectively. The symmetry of the structures after optimization was analyzed with FindSym [94] and Spglib [95], while momentum space analysis was performed within SeeK-path [96].

### v3

@@ IV Summary
Starting from the initial P6/mmm structure, which contains an ideal Fe kagome net, we found that there exists a critical value of the Hubbard parameter ( eV), for which the softened phonon mode becomes imaginary. From analysis of the most predominant imaginary soft branch at the L point, we found a new structure with the Ibam symmetry. This structure comes from P6/mmm due to the shift of Ge atoms within the Fe kagome net layer. However, the lattice dynamics of the Ibam structure reveals that it is also unstable with soft mode at the point. In this case, the displacement introduced by this soft mode associated with the shift of Ge atoms in a direction perpendicular to the Fe kagome net leads to a stable Immm structure. Condensation of both modes leads to the first-order structural phase transition, which is, in turn, responsible for the formation of the CDW. It is also worth noting that the occurrence of the first-order phase transition in FeGe is confirmed by experiments.
@@
For completeness, we performed also calculations for the slab structure with the Immm symmetry to generate STM images for the Fe-kagome-like and Ge-deformed honeycomb surface terminations. The calculated STM topographies indicate the existence of charge ordering. The observed CDW phase is associated with a deformed Ge honeycomb-like lattice.
Supplemental Material
Lattice dynamics study of electron-correlation-induced charge density wave
in antiferromagnetic kagome metal FeGe
Andrzej Ptok et al.
(Dated: April 28, 2025)
In this Supplemental Material, we present additional results:
Fig. S2 – Dependence of phonon dispersions in FeGe with P6/mmm and Immm symmetries calculated using different DFT+U variants: the simplified (rotationally invariant) approach to the DFT+U, introduced by Dudarev et al. [85] versus the rotationally invariant DFT+U introduced by Liechtenstein et al. [86].
Fig. S3 – Figure presenting the displacement of atoms induced by the discussed soft modes.
Fig. S4 – Comparison of the electronic band structure of the system with P6/mmm and Immm symmetries.
The electronic band structures for the P6/mmm and Immm magnetic unit cells are presented in Figs. S4(a) and S4(c), respectively. To facilitate comparison, we also present the folded band structure for the P6/mmm phase [Fig. S4(b)], calculated using a cell with a shape similar to the magnetic unit cell of the Immm phase [presented in Fig. 3(b)]. As we can see, inclusion of the SOC opens band gaps in several points of the Brillouin zone (see orange and blue lines in Fig. S4, to compare the electronic band structure in the absence and presence of the SOC). We can simply find similarities between the electronic band structure for Immm [Fig. S4 (c)] and the folded band structure for P6/mmm [see Fig. S4(b)].